Suppose that the marginal PDF of X is fX(X)=6x(1−x), for 0≤x ≤1, and theconditional PDF of Y given X =x is fY|X(y|x)= 1/ 1−x, for 0≤x ≤y ≤1. (a) Find E[Y |X =x]. (b) Find E[E[Y|X]].
Answer #1

f(x) = GX C1-x) 0604 f (y/2) = I LO Ely/x-x) = f y tay dy fag) = f(4/2). fQ 1.60 (1-x) = 60 Eylx=x) = f g 6x dy = 6x g ² e 3xECE MIX)] = E(Y) = jy fcyn dy - J y 3 gr dy = f 3y3dyPlease give thumbs up if you are satisfied else comment